OpenStudy (anonymous):

This question is associated with the first first law of thermodynamics. The question is as follows: A lead bullet initially at 30°C just melts upon striking a target. Assuming that all of the initial kinetic energy of the bullet goes into the internal energy of the bullet, calculate the impact speed of the bullet. Right. I've looked up the meltint temperature of lead, as well as the specific heat capacity and the latent heat of fusion. T_{melt} = 327K c = 0.13kJ/kg*K L_f = 23kJ/kg Considering all the kinetic energy is assumed to go into the internal energy I set up the foll

OpenStudy (anonymous):

$$Q = E_k \Rightarrow m(L_f+c(T_{melt}-T_0)) = \frac{1}{2}mv^2$$ where T_0 = 303K After simplifying and inserting the given numbers, I get the answer $$v = 11.1m/s$$, which according to the professor is extraordinarly wrong, as the correct answer should be $$v = 354m/s$$

OpenStudy (anonymous):

Obviously the melting point should be 327+273 = 600K